# Birational complexity and conic fibrations

### v1

@@ Abstract.
Let $(X,B)$ be a log Calabi–Yau pair of dimension $n$, index one, and birational complexity $c$. We show that $(X,B)$ has a crepant birational model that admits a tower of Mori fiber spaces of which at least $n-c$ are conic fibrations. Motivated by the proof of the previous statement, we introduce new measures of the complexity of a log Calabi–Yau pair; the alteration complexity and the conic complexity. We characterize when these invariants are zero. Finally, we give applications of the tools of the main theorem to birational superrigidity, Fano hypersurfaces, dual complexes, Weil indices of Fano varieties, and klt singularities.
@@ 1. Introduction
The complexity $c(X,B)$ of a log Calabi–Yau pair $(X,B)$ is defined to be $\dim X+\operatorname{rank}{\rm Cl}(X)_{\mathbb{Q}}-|B|$, where $|B|$ stands for the sum of the coefficients of $B$. The complexity gives an intrinsic measure of how far a given log Calabi–Yau pair is from being toric [11]. In [37], Svaldi and the author characterized toric singularities using a local variant of the complexity. In [21], Gongyo and the author introduced a variant of the complexity for generalized pairs. Using this variant, they proposed a conjecture that unifies the complexity, Mukai’s conjecture, and the Kobayashi–Ochiai Theorem. The aforementioned results lead to strong geometric consequences when the considered complexity measure is strictly smaller than one. However, not much is known when the complexity is larger than or equal to one.
In this article, we study the birational geometry of log Calabi–Yau pairs $(X,B)$ from the perspective of birational complexity. The birational complexity $c_{\rm bir}(X,B)$ is a birational variant of the complexity (see Definition 2.20). It was introduced by Mauri and the author as a proxy to study dual complexes [31]. In this work, we consider log Calabi–Yau pairs $(X,B)$ of index one and birational complexity $c$. We prove that some crepant birational model of $(X,B)$ admits a tower of Mori fiber spaces of which at least $\dim X-c$ are conic fibrations. In particular, this result sheds light on the birational geometry of $(X,B)$ when the birational complexity is smaller than the dimension.
@@ 1.1. Birational complexity and conic fibrations
The following theorem is the main result of this article.
@@
(c.f. Theorem 5.6) Let $(X,B)$ be a log Calabi–Yau pair of dimension $n$ and index one. Let $c:=c_{\rm bir}(X,B)$. Then $(X,B)$ has a crepant birational model that admits a tower of Mori fiber spaces of which at least $n-c$ are conic fibrations.
@@
Let $X_{d}$ be a smooth hypersurface of degree $d$ in $\mathbb{P}^{n}$. If $d\leq n$, then $X_{d}$ has a birational model that admits a tower of Mori fiber spaces of which at least $n-d$ are conic fibrations.
@@ 1.4. Applications
    In Theorem 5.10, we study the fundamental group of $U:=X\setminus B$ where $X$ is a Fano variety and $(X,B)$ is a $1$-complement [^1] A $1$-complement is a boundary $B$ for which $(X,B)$ is log Calabi–Yau and $K_{X}+B\sim 0$.. We show that $\pi_{1}^{\rm alg}(U)$ admits a normal abelian subgroup of index at most $c(n)$ and rank at most $n-k_{\rm alt}(X,B)$.
@@ 2. Preliminaries
[…]
@@ 2.1. Equivariant generalized pairs
such that $K_{X}+B+\mathbf{M}_{X}$ is $\mathbb{R}$-Cartier. We say that $(X,B,\mathbf{M})$ is a $G$-equivariant generalized pair if $B$ is an effective divisor. By [35, Proposition 2.4] the $G$-equivariant b-nef divisor $\mathbf{M}$ is the pull-back of a b-nef divisor $\mathbf{N}$ on the quotient $X/G$. If $Y$ is a birational model of $X/G$ on which $\mathbf{N}$ descends, then we say that $Y$ is a quotient by $G$ where $\mathbf{M}$ descends. Indeed, $Y$ is the quotient by $G$ of a $G$-equivariant birational model of $X$ (see [35, Proposition 2.4]). In the previous setting, we write $G\leqslant{\rm Aut}(X,B,\mathbf{M})$.
@@ 2.2. Fano type varieties and log Calabi–Yau pairs
Let $X\rightarrow W$ be a contraction. We say that $X$ is of Fano type over $W$ (or simply, Fano type over $W$) if there exists a boundary $B$ on $X$ that is big over $W$ and $(X,B)$ is a klt Calabi–Yau pair over $W$. If $X\rightarrow W$ is a Fano type morphism, then $X$ is a relative Mori dream space over $W$ (see [10, Theorem 3.18]). In particular, the MMP for any divisor on $X$ over $W$ terminates either with a good minimal model over $W$ or a Mori fiber space over $W$.
@@
Let $X\rightarrow W$ be a contraction. The variety $X$ is of Fano type over $W$ if and only there exists a boundary $B$ such that $(X,B)$ is klt and $-(K_{X}+B)$ is big and nef over $W$.
@@
By [37, Lemma 3.7], we may assume that the b-nef divisor $\mathbf{M}$ is trivial.
@@
The second statement is proved in [4, Lemma 2.12].
We prove the third statement. Let $(X,\Delta)$ be a klt pair for which $-(K_{X}+\Delta)$ is big and nef over $W$ provided by Lemma 2.16. Replacing $\Delta$ with $(1-\epsilon)B+\epsilon\Delta$ for $\epsilon>0$ small enough, we may assume that $Y\dashrightarrow X$ only extracts non-canonical places of $(X,\Delta)$. By [32, Theorem 1], there exists a projective birational morphism $Z\rightarrow X$ extracting exactly the exceptional divisors of $Y\dashrightarrow X$. Let $(Z,\Delta_{Z})$ be the log pull-back of $(X,\Delta)$ to $Z$. The pair $(Z,\Delta_{Z})$ is klt and $-(K_{Z}+\Delta_{Z})$ is big and nef over $W$. Hence, $Z$ is of Fano type over $W$ by Lemma 2.16. As $Y\dashrightarrow Z$ is small, we have that $Y\rightarrow W$ is of Fano type.
The fourth statement is proved in [34, Lemma 3.17].
We show the last statement. Let $Y\rightarrow X$ be a finite morphism such that its divisorial branch locus is contained in $\lfloor B\rfloor$. Let $Z\rightarrow X$ be the Galois closure of $Y\rightarrow X$. Hence, the divisorial branch locus of $Z\rightarrow X$ is contained in $\lfloor B\rfloor$. By Lemma 2.16, there is a klt pair $(X,\Delta)$ for which $-(K_{X}+\Delta)$ is big and nef over $W$. For each prime divisor $P$ on $X$, we let $m_{P}$ be the ramification index of $Z\rightarrow X$ at $P$. We replace $\Delta$ with $(1-\epsilon)B+\epsilon\Delta$ for $\epsilon>0$ small enough and assume that ${\rm coeff}_{P}(\Delta)>1-\frac{1}{m_{P}}$ for every $P$. Hence, the log pull-back $(Z,\Delta_{Z})$ of $(X,\Delta)$ to $Z$ is a log pair, i.e., $\Delta_{Z}$ is an effective divisor. Hence, the pair $(Z,\Delta_{Z})$ is klt and $-(K_{Z}+\Delta_{Z})$ is big and nef over $W$. We conclude that $Z$ is of Fano type over $W$ by Lemma 2.16. Hence, $Y$ is of Fano type over $W$ by the fourth statement. ∎
@@ 2.3. Complexity measures for log Calabi–Yau pairs
$\bar{c}(X/W,B,\mathbf{M}):=\inf\{c(\Sigma/W)\mid\text{ $\Sigma$ is a decomposition of $B$ }\}.$
In [37, Proposition 30], the authors show that the complexity is indeed a minimum, i.e., it is computed by some decomposition of $B$.
@@
$c_{\rm bir}(X/W,B,\mathbf{M}):={\rm inf}\left\{c(X^{\prime}/W,B^{\prime},\mathbf{M})\middle|\begin{gathered}\text{ $(X^{\prime},B^{\prime},\mathbf{M})$ is a crepant birational model}\\ \text{ of $(X,B,\mathbf{M})$ over $W$ and $B^{\prime}\geq 0$ }\end{gathered}\right\}.$
@@
$c_{\rm alt}(X/W,B,\mathbf{M}):={\rm inf}\left\{c_{\rm bir}(Y/W,B_{Y},\mathbf{N})\middle|\begin{gathered}\text{ $(Y,B_{Y},\mathbf{N})\rightarrow(X,B,\mathbf{M})$ is a crepant finite Galois}\\ \text{morphism over $W$ and $Y\rightarrow W$ is a contraction}\end{gathered}\right\}.$
@@
$k(X,B,\mathbf{M}):=\dim X-\max\{r(\mathcal{T})\mid\text{ $\mathcal{T}$ is a tower of fibrations for $X$ }\}.$
@@
$\min\{k_{\rm bir}(Y,B_{Y},\mathbf{N})\mid\text{ $(Y,B_{Y},\mathbf{N})\rightarrow(X,B,\mathbf{M})$ is a crepant finite Galois morphism }\}.$
@@
where $Y$ is a Bott $\mathbb{Q}$-tower variety of height $n-1$ and the $\mathcal{L}_{i}$’s are torus invariant $\mathbb{Q}$-line bundle on $Y$.
@@ 2.4. Negativity lemma
Since $Y$ is $\mathbb{Q}$-factorial, we can find an effective divisor $F$ on $X$ that is antiample over $Y$. Let $A$ be an ample divisor on $Y$. Then, the divisor $\phi^{*}(A)-\epsilon F$ is ample on $X$. Let $\delta>0$ be a positive rational number. Let $0\leq H\sim_{\mathbb{Q}}E+\delta\phi^{*}(A)-\delta\epsilon F$. Note that $H-E\sim_{\mathbb{Q}}\delta\epsilon F-\delta\phi^{*}(A)$ is anti-ample over $Y$. Furthermore, $\phi_{*}(H-E)=\phi_{*}H\geq 0$. By [27, Lemma 3.39], we conclude that $H\geq E$. ∎
@@
By contradiction, assume that a component $Q$ of $E$ is not contained in ${\rm Bs}_{-}(E)$. Let $p\colon Z\rightarrow X$ and $q\colon Z\rightarrow Y$ be a resolution of the birational map $\pi$. Let $A$ be an ample divisor on $X$. By assumption, for every $\epsilon>0$, the prime divisor $Q$ is not contained in ${\rm Bs}(E+\epsilon A)$. Hence, for every $\epsilon>0$, the effective divisor $p^{*}Q$ is not contained in ${\rm Bs}(p^{*}E+\epsilon p^{*}A)$. By [33, Lemma 1.27], we have ${\rm Bs}_{-}(p^{*}E)\subseteq\bigcup_{\epsilon>0}{\rm Bs}(p^{*}E+\epsilon p^{*}A)$. In particular, $p^{*}Q$ is not contained in ${\rm Bs}_{-}(p^{*}E)$. However, since $\pi$ is a birational contraction, the divisor $p^{*}E$ is exceptional over $Y$. Hence, by Lemma 2.27, we have $\operatorname{supp}p^{*}Q\subseteq\operatorname{supp}p^{*}E\subseteq{\rm Bs}_{-}(E)$. This leads to a contradiction. ∎

### v2

@@ Abstract.
Let $(X,B)$ be a log Calabi–Yau pair of dimension $n$, index one, and birational complexity $c$. We show that $(X,B)$ has a crepant birational model that admits a tower of Mori fiber spaces of which at least ncn-c are conic fibrations. The main tool to prove this statement is a geometric version of the Jordan property for finite actions on rationally connected varieties. Motivated by the proof of the previous statement, we introduce new measures of the complexity of a log Calabi–Yau pair; the alteration complexity and the conic complexity. We characterize when these invariants are zero. Finally, we give applications of the tools of the main theorem to birational superrigidity, Fano hypersurfaces, dual complexes, Weil indices of Fano varieties, and klt singularities.
@@ 1. Introduction
The complexity $c(X,B)$ of a log Calabi–Yau pair $(X,B)$ is defined to be $\dim X+\operatorname{rank}{\rm Cl}(X)_{\mathbb{Q}}-|B|$, where $|B|$ stands for the sum of the coefficients of $B$. The complexity gives an intrinsic measure of how far a given log Calabi–Yau pair is from being toric [11]. In [32], Svaldi and the author characterized toric singularities using a local variant of the complexity. In [21], Gongyo and the author introduced a variant of the complexity for generalized pairs. Using this variant, they proposed a conjecture that unifies the complexity, Mukai’s conjecture, and the Kobayashi–Ochiai Theorem. The aforementioned results lead to strong geometric consequences when the considered complexity measure is strictly smaller than one. However, not much is known when the complexity is larger than or equal to one. In this article, we study the birational geometry of log Calabi–Yau pairs $(X,B)$ from the perspective of birational complexity. The birational complexity $c_{\rm bir}(X,B)$ is a birational variant of the complexity (see Definition 2.20). It was introduced by Mauri and the author as a proxy to study dual complexes [31]. In this work, we consider log Calabi–Yau pairs $(X,B)$ of index one and birational complexity $c$. We prove that some crepant birational model of $(X,B)$ admits a tower of Mori fiber spaces of which at least $\dim X-c$ are conic fibrations. In particular, this result sheds light on the birational geometry of $(X,B)$ when the birational complexity is smaller than the dimension.
1.1. Finite actions on rationally connected varieties
The main motivation of this article is to study how the birational complexity of a log Calabi–Yau pair reflects in its (birational) geometry. However, most of the statements that are proved in the article are reduced to problems about finite actions on rationally connected varieties. The Jordan property for finite actions on rationally connected varieties states the following. Let $n$ be a positive integer, if $G$ is a finite group acting faithfully on a $n$-dimensional rationally connected variety, then there exists $A\leqslant G$ a normal abelian subgroup of rank at most $n$ and index at most $c(n)$. This theorem was proved by Prokhorov and Shramov [40] by using the boundedness of Fano varieties proved by Birkar [6]. In other words, finite groups acting on rationally connected varieties are almost abelian, with respect to the dimension of the variety. In previous works, the author has studied how the existence of such a group $A$ reflects on the geometry of the rationally connected variety. In [36], it is proved that if $A=\mathbb{Z}/m\mathbb{Z}\leqslant X$ for $X$ a Fano type surface with $m$ large enough, then in an $A$-equivariant birational model $X^{\prime}$ of $X$, we have $A<\mathbb{G}_{m}<{\rm Aut}(X^{\prime})$. In [37], it is proved that if $A=(\mathbb{Z}/m\mathbb{Z})^{n}\leqslant X$ for $X$ a Fano type variety of dimension $n$ and $m$ large enough (compared with $n$), then $X$ is rational and it is indeed a compactification of $\mathbb{G}_{m}^{n}$. In [38], it is proved that if $A=(\mathbb{Z}/m\mathbb{Z})^{k}\leqslant X$ for $X$ a Fano type variety of dimension $n$ and $m$ large enough (compared with $n$), then the coregularity of $X$ is at most nkn-k. The following theorem gives a more geometric explanation of the Jordan property for finite groups acting on rationally connected varieties.
Theorem 1.1 (Geometric Jordan property for finite actions on rationally connected varieties).
Let $n$ be a positive integer. There exists a constant $c(n)$, only depending on $n$, satisfying the following. Let $G$ be a finite group acting on a $n$-dimensional rationally connected variety $X$. Then, there exists a normal abelian subgroup $A\leqslant G$ of rank at most $n$ and index at most $c(n)$ satisfying the following. There exists an $A$-equivariant birational map $X\dashrightarrow X^{\prime}$ and an $A$-equivariant fibration $X^{\prime}\rightarrow\mathbb{P}^{k}$ that induces an action of $A$ on $\mathbb{P}^{k}$ and this action factors through $\mathbb{G}_{m}^{k}$.
In other words, in some birational model, the group $A$ comes from a torus action on the base of a fibration. It is unclear under which conditions the torus action of $\mathbb{P}^{k}$ can be lifted to a torus action of $X^{\prime}$.
@@ 1.2. Birational complexity and conic fibrations
The following theorem is the second main result of this article.
@@
(c.f. Theorem 5.6) Let $(X,B)$ be a log Calabi–Yau pair of dimension $n$ and index one. Let $c:=c_{\rm bir}(X,B)$. Then $(X,B)$ has a crepant birational model that admits a tower of Mori fiber spaces of which at least ncn-c are conic fibrations.
@@
Let $X_{d}$ be a smooth hypersurface of degree $d$ in $\mathbb{P}^{n}$. If $d\leq n$, then $X_{d}$ has a birational model that admits a tower of Mori fiber spaces of which at least ndn-d are conic fibrations.
@@ 1.5. Applications
    In Theorem 5.10, we study the fundamental group of $U:=X\setminus B$ where $X$ is a Fano variety and $(X,B)$ is a $1$-complement [^1] A 11-complement is a boundary $B$ for which $(X,B)$ is log Calabi–Yau and $K_{X}+B\sim 0$.. We show that $\pi_{1}^{\rm alg}(U)$ admits a normal abelian subgroup of index at most $c(n)$ and rank at most $n-k_{\rm alt}(X,B)$.
@@ 2. Preliminaries
${(X,B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{g}$${(X,B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm id}_{W}}$${W}$
@@ 2.1. Equivariant generalized pairs
such that $K_{X}+B+\mathbf{M}_{X}$ is $\mathbb{R}$-Cartier. We say that $(X,B,\mathbf{M})$ is a $G$-equivariant generalized pair if $B$ is an effective divisor. By [38, Proposition 2.4] the $G$-equivariant b-nef divisor $\mathbf{M}$ is the pull-back of a b-nef divisor $\mathbf{N}$ on the quotient $X/G$. If $Y$ is a birational model of $X/G$ on which $\mathbf{N}$ descends, then we say that $Y$ is a quotient by $G$ where $\mathbf{M}$ descends. Indeed, $Y$ is the quotient by $G$ of a $G$-equivariant birational model of $X$ (see [38, Proposition 2.4]). In the previous setting, we write $G\leqslant{\rm Aut}(X,B,\mathbf{M})$.
@@ 2.2. Fano type varieties and log Calabi–Yau pairs
Let $X\rightarrow W$ be a contraction. We say that $X$ is of Fano type over $W$ (or simply, Fano type over $W$) if there exists a boundary $B$ on $X$ that is big over $W$ and $(X,B)$ is a klt Calabi–Yau pair over $W$. If $X\rightarrow W$ is a Fano type morphism, then $X$ is a relative Mori dream space over $W$ (see [9, Theorem 3.18]). In particular, the MMP for any divisor on $X$ over $W$ terminates either with a good minimal model over $W$ or a Mori fiber space over $W$.
@@
Let $X\rightarrow W$ be a contraction. The variety $X$ is of Fano type over $W$ if and only there exists a boundary $B$ such that $(X,B)$ is klt and $(K_{X}+B)$ is big and nef over $W$.
@@
By [32, Lemma 3.7], we may assume that the b-nef divisor $\mathbf{M}$ is trivial.
@@
The second statement is proved in [5, Lemma 2.12].
We prove the third statement. Let $(X,\Delta)$ be a klt pair for which $(K_{X}+\Delta)$ is big and nef over $W$ provided by Lemma 2.16. Replacing $\Delta$ with $(1-\epsilon)B+\epsilon\Delta$ for $\epsilon>0$ small enough, we may assume that $Y\dashrightarrow X$ only extracts non-canonical places of $(X,\Delta)$. By [34, Theorem 1], there exists a projective birational morphism $Z\rightarrow X$ extracting exactly the exceptional divisors of $Y\dashrightarrow X$. Let $(Z,\Delta_{Z})$ be the log pull-back of $(X,\Delta)$ to $Z$. The pair $(Z,\Delta_{Z})$ is klt and $(K_{Z}+\Delta_{Z})$ is big and nef over $W$. Hence, $Z$ is of Fano type over $W$ by Lemma 2.16. As $Y\dashrightarrow Z$ is small, we have that $Y\rightarrow W$ is of Fano type.
The fourth statement is proved in [37, Lemma 3.17].
We show the last statement. Let $Y\rightarrow X$ be a finite morphism such that its divisorial branch locus is contained in $\lfloor B\rfloor$. Let $Z\rightarrow X$ be the Galois closure of $Y\rightarrow X$. Hence, the divisorial branch locus of $Z\rightarrow X$ is contained in $\lfloor B\rfloor$. By Lemma 2.16, there is a klt pair $(X,\Delta)$ for which $(K_{X}+\Delta)$ is big and nef over $W$. For each prime divisor $P$ on $X$, we let $m_{P}$ be the ramification index of $Z\rightarrow X$ at $P$. We replace $\Delta$ with $(1-\epsilon)B+\epsilon\Delta$ for $\epsilon>0$ small enough and assume that ${\rm coeff}_{P}(\Delta)>1-\frac{1}{m_{P}}$ for every $P$. Hence, the log pull-back $(Z,\Delta_{Z})$ of $(X,\Delta)$ to $Z$ is a log pair, i.e., $\Delta_{Z}$ is an effective divisor. Hence, the pair $(Z,\Delta_{Z})$ is klt and $(K_{Z}+\Delta_{Z})$ is big and nef over $W$. We conclude that $Z$ is of Fano type over $W$ by Lemma 2.16. Hence, $Y$ is of Fano type over $W$ by the fourth statement. ∎
@@ 2.3. Complexity measures for log Calabi–Yau pairs
[…]
In [32, Proposition 30], the authors show that the complexity is indeed a minimum, i.e., it is computed by some decomposition of $B$.
@@
[…]
@@
[…]
@@
${\mathcal{T}:X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}$${X_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{2}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{r}}$${X_{r}}$
@@
[…]
@@
where $Y$ is a Bott $\mathbb{Q}$-tower variety of height n1n-1 and the $\mathcal{L}_{i}$’s are torus invariant $\mathbb{Q}$-line bundle on $Y$.
@@ 2.4. Negativity lemma
Since $Y$ is $\mathbb{Q}$-factorial, we can find an effective divisor $F$ on $X$ that is antiample over $Y$. Let $A$ be an ample divisor on $Y$. Then, the divisor $\phi^{*}(A)-\epsilon F$ is ample on $X$. Let $\delta>0$ be a positive rational number. Let $0\leq H\sim_{\mathbb{Q}}E+\delta\phi^{*}(A)-\delta\epsilon F$. Note that $H-E\sim_{\mathbb{Q}}\delta\epsilon F-\delta\phi^{*}(A)$ is anti-ample over $Y$. Furthermore, $\phi_{*}(H-E)=\phi_{*}H\geq 0$. By [25, Lemma 3.39], we conclude that $H\geq E$. ∎
@@
By contradiction, assume that a component $Q$ of $E$ is not contained in ${\rm Bs}_{-}(E)$. Let $p\colon Z\rightarrow X$ and $q\colon Z\rightarrow Y$ be a resolution of the birational map $\pi$. Let $A$ be an ample divisor on $X$. By assumption, for every $\epsilon>0$, the prime divisor $Q$ is not contained in ${\rm Bs}(E+\epsilon A)$. Hence, for every $\epsilon>0$, the effective divisor $p^{*}Q$ is not contained in ${\rm Bs}(p^{*}E+\epsilon p^{*}A)$. By [35, Lemma 1.27], we have ${\rm Bs}_{-}(p^{*}E)\subseteq\bigcup_{\epsilon>0}{\rm Bs}(p^{*}E+\epsilon p^{*}A)$. In particular, $p^{*}Q$ is not contained in ${\rm Bs}_{-}(p^{*}E)$. However, since $\pi$ is a birational contraction, the divisor $p^{*}E$ is exceptional over $Y$. Hence, by Lemma 2.27, we have $\operatorname{supp}p^{*}Q\subseteq\operatorname{supp}p^{*}E\subseteq{\rm Bs}_{-}(E)$. This leads to a contradiction. ∎